Suppose now that country's national income increases to $13 billion. Assuming the amount paid in taxes is fixed at $3 billion and MPC = 0.75, what will be the new household consumption?



1
Expert's answer
2022-04-06T14:06:13-0400

Solution

"The multiplier is \frac{1}{ (1 - MPC)}"

"= \frac{1}{MPS}"


"= \frac{1 }{0.25}"

= 4.